OpenStudy (anonymous):

All edges of a cube are expanding at a rate of 6 cm/sec. How fast is the volume changing when each edge is (a) 2 cm and (b) 10 cm?

OpenStudy (anonymous):

a lovely related rate problem for sure.

OpenStudy (anonymous):

\[V=x^3\] so you know \[V'=3x^2x'\] and you are told that \[x'=6\] so you have \[V'=18x^2\] now replace x by your two values to get the answer

OpenStudy (anonymous):

im confused....how did u get v'.

OpenStudy (anonymous):

i just wrote it. in other words we know \[V(x(t))=x^3(t)\]

OpenStudy (anonymous):

arent we suppsed to differentiate....liek write dr/dt = smthng adn then dv/dt = smthng? thats how we didi it in class but idk how to do it 4 spheres

OpenStudy (anonymous):

we can write it like that if you like. instead of writing \[\frac{dx}{dt}\] i just wrote \[x'\]

OpenStudy (anonymous):

you can write \[V=x^3\] \[\frac{dV}{dt}=3x^2\frac{dx}{dt}\] if you prefer

OpenStudy (anonymous):

still confused?

OpenStudy (anonymous):

maybe in class you wrote \[\frac{dV}{dt}=\frac{dV}{dx}\times \frac{dx}{dt}\]

OpenStudy (anonymous):

i understand it this way...wat are teh rest of teh steps wen we wrtie it this way in terms of dv//dt...

OpenStudy (anonymous):

just as i wrote it. you are told that \[\frac{dx}{dt}=6\] in the statement "All edges of a cube are expanding at a rate of 6 cm/sec."

OpenStudy (anonymous):

and you know if \[V=x^3\] then \[\frac{dV}{dx}=3x^2\]

OpenStudy (anonymous):

put that together and you get \[\frac{dV}{dt}=\frac{dV}{dx}\times \frac{dx}{dt}=3x^2\times 6=18x^2\]

OpenStudy (anonymous):

so now we plug 2 and 10 into it to get the ansewrs?

OpenStudy (anonymous):

there is a second part to teh question....can u help?

OpenStudy (anonymous):

determine how fast the surface area is changing when each edge is a) 2 and b) 10 cm.

OpenStudy (anonymous):

yes you plug in the numbers to get the answer

OpenStudy (anonymous):

surface area of a cube is \[S=6x^2\] since there are six sides each with area \[x^2\] repeat the process and get \[S'=12xx'\] or \[S'=72x\]

OpenStudy (anonymous):

since u explained this well may u help me w/ another one? this one is realll confusing me b.c it is a cone

OpenStudy (anonymous):

At a sand and gravel plant, sand is falling off a conveyor and onto a conical pile at a rate of 10 cubic ft/ min. The diameter of the base of the cone is approximately 3x the altitude. At what rate is the height of the pile changing when the pile is 15 ft high?

OpenStudy (anonymous):

so what formula do we use to plug in 2 and 10? is it V=3x^2dx/dt?